A rectangle has an area of x^3 + 5x^2 + 5x – 2 square meters and a width of x + 2 meters. Find its length.
Answer by ewatrrr(24785)

 
Hi
A = Lw
L = ( x^3 + 5x^2 + 5x – 2)/(x+2)
........
Using Synthetic Division
-2 1 5 5 -2
-2 -6 2
1 3 -1 0
A = (x+2)(x^2 + 3x - 1)
L = (x^2 + 3x - 1)